What is the effect on the area of a triangle if the base is doubled and the height is cut in half? What happens to the area, if the base is doubled and the height remains the same?
Answer by stanbon(75887)

Area = (1/2)base*h
--------------------
What is the effect on the area of a triangle if the base is doubled and the height is cut in half?
Area = (1/2)(2b)(h/2) = (1/2)bh
No change in Area
----------------------------
What happens to the area, if the base is doubled and the height remains the same?
Area = (1/2)(2b)h = 2[(1/2)bh]
The area is doubled.
=========================
Cheers,
Stan H.